• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Armstrong

    Log entry time 00:00:34 on July31,2003

    Entry number 108226

    keyword=Swing Shift Summary


    Shift Log   Swing shift   July 30 2003


    D.S. Armstrong




    Shift Start:

    holding field: 0°
    target HWP: out
    lasers: 4,6,7 on
    Target: Pol 3He
    Cell: Priapus
    Beam: waiting for beam (linac energy change)



    Septum continues to cool.
    HRS magnets at new setpoints.

    18:34    Start Target Temperature study (as instructed
    by Vince).

    Target Temperature study: See HaLog entry 108222.

    22:30    Septum coils each below 7K. Start Ramping
    up current. Ask for 135 A first.

    23:20    Septum reached 135 A; we sat there for 20 minutes,
    coil temperatures dropped nicely. Now ask for setpoint of 270.6 A.

    23:50    Septum at 270 A.

    Shift End:


    holding field: 0°
    target HWP: out
    lasers: longitudinal: 4,6,7 ON
    Target: 3He cell
    Cell: Priapus
    Beam: waiting for beam



    the babysitter hands off to owl shift, and goes home to sleep...